# Design of low-energy transfers in cislunar space using sequences of lobe dynamics

### v1

@@ Abstract
Dynamical structures in the circular restricted three-body problem (CR3BP) are fundamental for designing low-energy transfers, as they aid in analyzing phase space transport and designing desirable trajectories. This study focuses on lobe dynamics to exploit local chaotic transport around celestial bodies, and proposes a new method for systematically designing low-energy transfers by combining multiple lobe dynamics. A graph-based framework is constructed to explore possible transfer paths between departure and arrival orbits, reducing the complexity of the combinatorial optimization problem for designing fuel-efficient transfers. Based on this graph, low-energy transfer trajectories are constructed by connecting chaotic orbits within lobes. The resulting optimal trajectory in the Earth–Moon CR3BP is then converted into an optimal transfer in the bicircular restricted four-body problem using multiple shooting. The obtained transfer is compared with existing optimal solutions to demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Trajectory design methods in literature fall into two main categories: numerical and geometrical. Numerical approaches (e.g., Refs. [88, 98, 71, 30]) identify transfer trajectories by solving optimization problems. By evaluating a large set of initial conditions, these methods can find a wide range of solutions for fuel consumption and time of flight, often yielding Pareto-optimal sets. Their drawbacks include the need for accurate initial guesses and high computational cost. Geometrical or dynamical systems approaches (e.g., Refs. [48, 50, 29, 62, 35, 37, 89]) use the natural dynamics described by dynamical systems theory to generate low-energy transfers. These approaches are limited to planar problems and often result in longer transfer times. The Genesis mission [56, 11] was the first to demonstrate the practical value of this approach by designing its nominal trajectory using modern dynamical systems theory. Both approaches are valuable and complementary, but further development of the geometrical approach is crucial to address the key challenge in trajectory design. The geometrical approach identifies desirable trajectories by analyzing phase space transport, while the numerical approach focuses on local dynamics near given points. Although the numerical approach may imply underlying dynamical structures, the geometrical approach is generally more effective at providing a variety of trajectories with specific properties.
The geometrical approach produces low-energy transfers based on chaotic transport in natural flows, as described by the theories of lobe dynamics and tube dynamics (e.g., Refs. [47, 51, 82]). These theories utilize invariant manifolds associated with periodic orbits in the planar circular restricted three-body problem (CR3BP) [96]. The four-dimensional phase space in the planar CR3BP is often analyzed using a two-dimensional Poincaré map. If periodic orbits appear as points on a Poincaré map, their manifolds may reveal lobe dynamics; otherwise, the manifolds may describe tube dynamics. Thus, it is important to analyze the manifolds of periodic orbits with an appropriate Poincaré map.
@@
Tube dynamics [48, 60] describes global transport between two celestial bodies based on the stable and unstable manifolds of libration point orbits [47], especially around $L_{1}$ or $L_{2}$ in the planar CR3BP. These manifolds are called Conley–McGehee tubes or simply tubes [60]. The tubes around $L_{1}$ and $L_{2}$ provide a framework for understanding the dynamics in the planar CR3BP, as shown in Ref. [50]. For example, the natural orbit transitions of Jupiter comets are described using tube dynamics in the Sun–Jupiter system [49]. Tube dynamics also explains the mechanism of ballistic lunar transfers through Earth–Moon $L_{2}$, enabling systematic design of such transfers [46]. One limitation of tube dynamics is that tubes associated with $L_{1}$ Lyapunov orbits do not help depart from the primary, as they move away from it in the CR3BP. Another limitation is the difficulty of applying this method to interplanetary transfers in the Solar System [82]. For example, tubes in the Sun–Earth and Sun–Mars planar CR3BP do not naturally intersect [82, 66].
@@
Lobe dynamics describes finer transport between regions around a single celestial body [47]. This theory was first developed to analyze fluid mixing in two-dimensional inviscid incompressible flows [83] and then extended to two-dimensional, area- and orientation-preserving maps [84]. In such maps, segments of the stable and unstable manifolds associated with periodic points may form small enclosed regions called lobes, which are key to transport by lobe dynamics. Within the theory of transport in dynamical systems [63, 64], this idea has been used to analyze the chaotic transport of phase space volume between regions in various fields, such as fluid dynamics [12, 24, 59, 108, 94], geophysical flows [109, 81, 23], celestial mechanics [82, 72], and chemical reaction dynamics [53, 44]. Computational tools for lobe area and transport rate [20, 68] have been developed, since applying lobe dynamics is limited by the computational difficulty of numerically identifying lobes [47]. In spacecraft trajectory design, lobe dynamics typically leverages the stable and unstable manifolds of resonant orbits. Previous studies [86, 31] have developed trajectory design methods that use multiple gravity assists of Jupiter’s moons by numerically exploiting lobe dynamics in the Jovian system. Further, Refs. [69, 70] have used the lobe dynamics of the 33:11 resonant orbit to design transfers to the Moon in the Earth–Moon system. In summary, lobe dynamics has been numerically investigated for moon-to-moon transfers or analytically applied around a single resonant orbit for trajectory design in the literature. These studies have yet to provide a simple interpretation of phase space transport for trajectory design based on lobe dynamics.
@@
Unlike the previous studies [86, 31, 69, 70], this study proposes a new method that systematically incorporates the dynamical structures of lobes associated with multiple resonant orbits into trajectory design. Earlier work by the authors’ group [37] provided the first theoretical investigation into the combination of multiple lobe dynamics in the standard map [17] via a graph-based framework. That Letter [37] also included a preliminary application to the planar CR3BP. This paper fully develops and validates a trajectory design method based on this approach. The proposed method is applied to construct Earth–Moon transfers in both the CR3BP and the bicircular restricted four-body problem (BCR4BP) [40, 10]. Numerical results demonstrate that lobe dynamics can reveal connections between a celestial body and a tube, which is difficult to achieve using tube dynamics alone. The optimal transfer obtained in the BCR4BP is compared with existing solutions to demonstrate the effectiveness of the proposed method.
The remainder of this paper is organized as follows: Section 2 describes the dynamical model of a spacecraft and related fundamental concepts. Section 3 explains lobe dynamics in detail, focusing on its application to the planar CR3BP. Section 4 presents a new method leveraging lobe dynamics for transfer trajectory design, based on Ref. [37]. Selected lobes are connected with appropriate impulsive inputs, and the overall transfer structure is summarized as a graph representing possible transfer paths. This graph enables identification of the optimal transfer connecting the start and goal orbits. Section 5 applies the proposed method to construct fuel-optimal Earth–Moon transfers. Finally, Section 6 discusses the role of lobe dynamics and the extension of the proposed method to the BCR4BP.
@@ 2 Preliminaries
This section introduces fundamental knowledge related to this study. The Earth–Moon CR3BP is mainly employed in this paper, while the Sun–Earth–Moon BCR4BP is used in Section 6 to compare the obtained results with those in literature. Note that lobe dynamics is described separately in the next section.
@@ 2.1 Circular restricted three-body problem (CR3BP)
This subsection describes key features of the CR3BP, whose coordinate system is shown in Fig. 1. This problem considers two celestial bodies, $P_{1}$ and $P_{2}$, and a spacecraft $P_{3}$, all treated as point masses. Let $m_{1}$, $m_{2}$ and $m_{3}$ denote the masses of $P_{1}$, $P_{2}$ and $P_{3}$, respectively, with the assumption that $m_{1}>m_{2}\gg m_{3}$. This condition means $P_{3}$ has no effect on the motion of $P_{1}$ and $P_{2}$. Another assumption in the CR3BP is that $P_{1}$ and $P_{2}$ move in a circular orbit around their barycenter. Under these assumptions, the equations of motion of $P_{3}$ are expressed in a non-dimensional form with a pseudo-potential function $U$ as follows:
@@
The CR3BP has several characteristics useful for trajectory design. First, the CR3BP has five libration points satisfying $\nabla U=\bm{0}$, denoted as $L_{i}$ ($i=1,2,\cdots,5$). The dynamics around these points has been well investigated in literature (e.g., Refs. [22, 78, 16]) because each libration point $L_{i}$ has families of periodic orbits. These periodic orbits and their stable and unstable manifolds are useful for analyzing the dynamics and designing trajectories. In addition, this model possesses one integral of motion, called the Jacobi constant $C_{J}$, which is defined as
@@ 3.2 Lobe sequences in the CR3BP
Figure 9 demonstrates effective lobes when $r_{L}^{\ast}=0.002$ in the selected lobe sequences. In Fig. 9, the name of a lobe sequence, LX.Y, indicates the Y-th lobe sequence of the X-th periodic orbit. Here, the first orbit is the 77:22 unstable resonant orbit, and the second orbit is the 33:11 unstable resonant orbit. The number in Fig. 9 shows the order of transfers in the lobe sequence.
@@ 4 Trajectory design based on lobe dynamics
This section describes the proposed method for designing low-energy chaotic transfers based on lobe dynamics. Effective lobe sequences are utilized to predict the motion of chaotic orbits within lobes. The selected chaotic orbits are connected by impulsive maneuvers to construct a low-energy transfer. These maneuvers help to overcome the partial barriers [64] formed by cantori [57] or the boundaries of resonance regions in the periapsis Poincaré map of the planar CR3BP. In the following sections, the threshold radius for effective lobes is fixed at $r_{L}^{\ast}=0.002$ as a representative example.
@@ 4.1 Overview of the design method
where $J$ is the sum of $w_{k}$ along a transfer path from one of the start nodes to one of the goal nodes, and $w^{\ast}$ is a weight threshold. The number of the combinations is reduced by only considering edges whose weight satisfies Eq. (23). The second constraint of Eq. (24) implies that if a spacecraft goes through one of the lobe sequences, at least two adjacent lobes in the lobe sequence must be included in the transfer path (see Fig. 12). This constraint requires the full exploitation of lobe dynamics during transfers, which demonstrates the effectiveness of the approach. Note that this condition can be neglected for mission design.
@@ 4.2 Test problem settings
A test problem for the fuel-optimal transfer is considered to investigate the design of each transfer arc and its corresponding weight. In this problem, the departure orbit and destination are defined within the periapsis Poincaré map, as summarized in Fig. 13. The departure orbit is the 77:22 stable resonant orbit in Fig. 14, and the triangles in Figs. 13 and 14 indicate its periapses. It is noteworthy that the altitudes of these periapses to the Earth are from $27,279$ km to $37,746$ km, implying that the starting points exist roughly around the geosynchronous Earth orbit (GEO) [105] since the altitude of the GEO is about $35,780$ km. The destination is defined as the left half of the stable manifold of the $L_{1}$ Lyapunov orbit, as shown in Fig. 15. The black line is the $L_{1}$ Lyapunov orbit, and the green lines represent its stable manifold. The blue regions in Figs. 13 and 15 illustrate the first intersection to the periapsis Poincaré map. These blue regions can be regarded as the exit to the moon realm, as a spacecraft within them is delivered to the Moon or possibly to deep space via tube dynamics. The other colored regions in Fig. 13 represent the selected effective lobe sequences corresponding to those in Fig. 9. Thus, this test problem aims to design a low-energy, fuel-optimal transfer for escaping from the Earth.
@@
4.3 Test results
Test results for all paths in both cases are summarized in Table 1 and Fig. 18. The effect of the constraint of Eq. (24) is also examined. The label “Estimated” indicates that the total costs are estimated by assuming that the edge weights for transfers based on lobe dynamics are precisely zero. The label “Without targeting” denotes that the total costs are determined by the graph in Case 1. The last label “With targeting” represents that the total costs are calculated from the graph in Case 2. Table 1 indicates the number of transfers found by the three approaches. In Case 1 (without targeting), some transfer paths do not yield the corresponding transfer trajectories because controlled transfer arcs fail to be reconstructed as a result of the propagation. In contrast, in Case 2 (with targeting), all transfer paths generate transfer trajectories due to the targeting strategy. The constraint of Eq. (24) reduces the number of transfers found, regardless of the difference between the approaches. Next, Fig. 18 shows swarm charts of the total transfer costs determined by the three approaches. A swarm chart displays each data point individually, clearly illustrating the distribution of discrete data. Figures 18 and 18 highlight low-cost transfers in Figs. 18 and 18, respectively. Figure 18 illustrates that transfer costs in Case 1 (without targeting) are significantly different from the estimated values. In Case 2 (with targeting), transfer costs have a distribution similar to the estimated values. Thus, the targeting strategy is practical for realizing transfer trajectories from the graph structure in the proposed method. The comparison between Fig. 18 and Fig. 18, and between Fig. 18 and Fig. 18 reveals that the number of transfers found is reduced by the constraint of Eq. (24), but their distribution does not change significantly.
@@ 4.3 Test results
Figure 18: Swarm chart for the transfer costs of all potential paths with different approaches.

### v2

@@ Abstract
Dynamical structures in the circular restricted three-body problem (CR3BP) are fundamental for designing low-energy transfers, as they aid in analyzing phase space transport and designing desirable trajectories. One of these dynamical structures, lobe dynamics, can be exploited to achieve local chaotic transport around celestial bodies. This study proposes and fully validates a systematic method for designing low-energy transfers by combining multiple sequences of lobe dynamics, building upon our prior preliminary framework. A graph-based framework is developed to explore possible transfer paths between departure and arrival orbits, reducing the complexity of the combinatorial optimization problem for fuel-efficient transfer design. Using this graph, low-energy transfer trajectories are constructed by connecting chaotic orbits within lobes. The resulting optimal trajectory in the Earth–Moon CR3BP is then converted into an optimal transfer in the bicircular restricted four-body problem via multiple shooting. This transfer is compared with existing optimal solutions to demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Trajectory design methods in the literature fall into two main categories: numerical and geometrical. Numerical approaches (e.g., Refs. [88, 98, 71, 30]) identify transfer trajectories by solving optimization problems. By evaluating a large set of initial conditions, these methods can find a wide range of solutions for fuel consumption and time of flight, often yielding Pareto-optimal sets. Their drawbacks include the need for accurate initial guesses and high computational cost. Geometrical or dynamical systems approaches (e.g., Refs. [48, 50, 29, 62, 35, 37, 89]) use the natural dynamics described by dynamical systems theory to generate low-energy transfers. These approaches are limited to planar problems and often result in longer transfer times. The Genesis mission [56, 11] was the first to demonstrate the practical value of this approach by designing its nominal trajectory using modern dynamical systems theory. Both approaches are valuable and complementary, but further development of the geometrical approach is crucial to address the key challenge in trajectory design. The geometrical approach identifies desirable trajectories by analyzing phase space transport, while the numerical approach focuses on local dynamics near given points. Although the numerical approach may imply underlying dynamical structures, the geometrical approach is generally more effective at providing a variety of trajectories with specific properties.
The geometrical approach produces low-energy transfers by exploiting chaotic transport in natural flows, which is described by the theories of tube dynamics and lobe dynamics (e.g., Refs. [47, 51, 82]). Both theories characterize the transport structures formed by invariant manifolds associated with periodic orbits in the planar circular restricted three-body problem (CR3BP) [96], but from different perspectives. Tube dynamics describes transport as a continuous flow along invariant manifolds, whereas lobe dynamics represents it as discrete transfer across manifold intersections. In astrodynamics, these two transport theories can be applied to interpret dynamics at different scales and in different regions, as detailed below.
@@
Tube dynamics [48, 60] describes global transport between two celestial bodies based on the stable and unstable manifolds of libration point orbits [47], especially around $L_{1}$ or $L_{2}$ in the planar CR3BP. These manifolds are called Conley–McGehee tubes or simply tubes [60]. The tubes around $L_{1}$ and $L_{2}$ provide a framework for understanding the dynamics in the planar CR3BP, as shown in Ref. [50]. For example, the natural orbit transitions of Jupiter comets are described using tube dynamics in the Sun–Jupiter system [49]. Tube dynamics also explains the mechanism of ballistic lunar transfers through Earth–Moon $L_{2}$ and enables their systematic design [46]. One limitation of tube dynamics is that tubes associated with $L_{1}$ Lyapunov orbits do not facilitate departure from the primary, as they move away from it in the CR3BP. Another limitation is its limited applicability to interplanetary transfers in the Solar System [82]. For example, tubes in the Sun–Earth and Sun–Mars planar CR3BP do not naturally intersect [82, 66].
@@
Lobe dynamics describes finer transport between regions around a single celestial body [47]. This mechanism complements tube dynamics by revealing pathways that connect the vicinity of the primary to tubes, which is difficult to achieve using tube dynamics alone. The theory of lobe dynamics was first developed to analyze fluid mixing in two-dimensional inviscid incompressible flows [83] and then extended to two-dimensional, area- and orientation-preserving maps [84]. In such maps, segments of the stable and unstable manifolds associated with periodic points may form small enclosed regions called lobes, which are key elements of transport by lobe dynamics. Within the theory of transport in dynamical systems [63, 64], this idea has been used to analyze the chaotic transport of phase space volume between regions in various fields, such as fluid dynamics [12, 24, 59, 108, 94], geophysical flows [109, 81, 23], celestial mechanics [82, 72], and chemical reaction dynamics [53, 44]. Computational tools for lobe area and transport rate [20, 68] have been developed, as the application of lobe dynamics is limited by the computational difficulty of identifying lobes numerically [47]. In spacecraft trajectory design, lobe dynamics typically leverages the stable and unstable manifolds of resonant orbits. Previous studies [86, 31] have developed trajectory design methods using multiple gravity assists of Jupiter’s moons by numerically exploiting lobe dynamics in the Jovian system. Furthermore, Refs. [69, 70] have used the lobe dynamics of the 33:11 resonant orbit to design transfers to the Moon in the Earth–Moon system. In summary, lobe dynamics has been investigated numerically for moon-to-moon transfers or analytically for trajectory design around a single resonant orbit in the literature. These studies have not yet provided a simple interpretation of phase space transport for trajectory design based on lobe dynamics.
@@
Unlike the previous studies [86, 31, 69, 70], this study proposes a new method that systematically incorporates the dynamical structures of lobes associated with multiple resonant orbits into trajectory design. Earlier work by the authors’ group [37] provided the first theoretical investigation into the combination of multiple lobe dynamics in the standard map [17] via a graph-based framework. That Letter [37] also included a preliminary application to the planar CR3BP. This paper fully develops and validates a trajectory design method based on this approach. The proposed method is applied to construct Earth–Moon transfers in both the CR3BP and the bicircular restricted four-body problem (BCR4BP) [40, 10]. The optimal transfer obtained in the BCR4BP is compared with existing solutions to demonstrate the effectiveness of the proposed method.
@@ 2 Preliminaries
This section introduces fundamental knowledge related to this study. The Earth–Moon CR3BP is mainly employed in this paper, while the Sun–Earth–Moon BCR4BP is used in Section 6 to compare the results with those in the literature. Note that lobe dynamics is described separately in the next section.
@@ 2.1 Circular restricted three-body problem (CR3BP)
This subsection describes key features of the CR3BP, whose coordinate system is shown in Fig. 1. This problem considers two celestial bodies, $P_{1}$ and $P_{2}$, and a spacecraft $P_{3}$, all treated as point masses. Let $m_{1}$, $m_{2}$ and $m_{3}$ denote the masses of $P_{1}$, $P_{2}$ and $P_{3}$, respectively, with the assumption that $m_{1}>m_{2}\gg m_{3}$. This condition means $P_{3}$ has no effect on the motion of $P_{1}$ and $P_{2}$. Another assumption in the CR3BP is that $P_{1}$ and $P_{2}$ move in circular orbits around their barycenter. Under these assumptions, the equations of motion of $P_{3}$ are expressed in a non-dimensional form with a pseudo-potential function $U$ as follows:
@@
The CR3BP has several characteristics useful for trajectory design. First, the CR3BP has five libration points satisfying $\nabla U=\bm{0}$, denoted as $L_{i}$ ($i=1,2,\cdots,5$). The dynamics around these points have been well investigated in the literature (e.g., Refs. [22, 78, 16]) because each libration point $L_{i}$ has families of periodic orbits. These periodic orbits and their stable and unstable manifolds are useful for analyzing the dynamics and designing trajectories. In addition, this model possesses one integral of motion, called the Jacobi constant $C_{J}$, which is defined as
@@ 3.2 Lobe sequences in the CR3BP
Figure 9: Histogram of lobe radii for each lobe sequence.
Figure 9 illustrates the distribution of lobe radii in each lobe sequence. The name of a lobe sequence, LX.Y, indicates the Y-th lobe sequence corresponding to the X-th periodic orbit. Specifically, the first and second orbits are the 77:22 and 33:11 unstable resonant orbits, respectively. Note that lobes with $r_{L}^{\ast}<0.001$ are not thoroughly investigated because of numerical difficulties in detecting finer lobes. Figure 10 shows the effective lobes when $r_{L}^{\ast}=0.002$ for the selected lobe sequences. The numbers in this figure indicate the order of transfers in the lobe sequence.
@@ 4 Trajectory design based on lobe dynamics
This section describes the proposed method for designing low-energy chaotic transfers based on lobe dynamics. Effective lobe sequences are utilized to predict the motion of chaotic orbits within lobes. The selected chaotic orbits are connected by impulsive maneuvers to construct a low-energy transfer. These maneuvers help to overcome the partial barriers [64] formed by cantori [57] or the boundaries of resonance regions in the periapsis Poincaré map of the planar CR3BP. In the following sections except Section 4.4, the threshold radius for effective lobes is fixed at $r_{L}^{\ast}=0.002$ as a representative example.
@@ 4.1 Overview of the design method
where $J$ is the sum of $w_{k}$ along a transfer path from one of the start nodes to one of the goal nodes, and $w^{\ast}$ is a threshold weight. The number of the combinations is reduced by considering only edges whose weight satisfies Eq. (23). The second constraint of Eq. (24) implies that if a spacecraft goes through one of the lobe sequences, at least two adjacent lobes in the lobe sequence must be included in the transfer path (see Fig. 13). This constraint requires the full exploitation of lobe dynamics during transfers, which demonstrates the effectiveness of the approach. Note that this condition can be neglected for mission design.
@@ 4.2 Test problem settings
A test problem for the fuel-optimal transfer is considered to investigate the design of each transfer arc and its corresponding weight. In this problem, the departure orbit and destination are defined within the periapsis Poincaré map, as summarized in Fig. 14. The departure orbit is the 77:22 stable resonant orbit in Fig. 15, and the triangles in Figs. 14 and 15 indicate its periapses. It is noteworthy that the altitudes of these periapses to the Earth are from $27,279$ km to $37,746$ km, implying that the starting points exist roughly around the geosynchronous Earth orbit (GEO) [105] since the altitude of the GEO is about $35,780$ km. The destination is defined as the left half of the stable manifold of the $L_{1}$ Lyapunov orbit, as shown in Fig. 16. The black line is the $L_{1}$ Lyapunov orbit, and the green lines represent its stable manifold. The blue regions in Figs. 14 and 16 illustrate the first intersection to the periapsis Poincaré map. These blue regions can be regarded as the exit to the moon realm, as a spacecraft within them is delivered to the Moon or possibly to deep space via tube dynamics. The other colored regions in Fig. 14 represent the selected effective lobe sequences corresponding to those in Fig. 10. Thus, this test problem aims to design a low-energy, fuel-optimal transfer for escaping from the Earth.
@@
4.3 Test-problem results
@@
Figure 19: Swarm charts for the transfer costs of all potential paths with different approaches.
Test results for all paths in both cases are summarized in Table 1 and Fig. 19. The effect of the constraint of Eq. (24) is also examined. The label “Estimated” indicates that the total costs are estimated by assuming that the edge weights for transfers based on lobe dynamics are precisely zero. The label “Without targeting” denotes that the total costs are determined by the graph in Case 1. The last label “With targeting” represents that the total costs are calculated from the graph in Case 2. Table 1 indicates the number of transfers found by the three approaches. In Case 1 (without targeting), some transfer paths do not yield the corresponding transfer trajectories because controlled transfer arcs fail to be reconstructed as a result of the propagation. In contrast, in Case 2 (with targeting), all transfer paths generate transfer trajectories due to the targeting strategy. The constraint of Eq. (24) reduces the number of transfers found, regardless of the difference between the approaches. Next, Fig. 19 shows swarm charts of the total transfer costs determined by the three approaches. A swarm chart displays each data point individually, clearly illustrating the distribution of discrete data. Figures 19 and 19 highlight low-cost transfers in Figs. 19 and 19, respectively. Figure 19 illustrates that transfer costs in Case 1 (without targeting) are significantly different from the estimated values. In Case 2 (with targeting), transfer costs have a distribution similar to the estimated values. Thus, the targeting strategy is practical for realizing transfer trajectories from the graph structure in the proposed method. The comparison between Fig. 19 and Fig. 19, and between Fig. 19 and Fig. 19 reveals that the number of transfers found is reduced by the constraint of Eq. (24), but their distribution does not change significantly.
@@
4.4 Impact of the threshold radius $r_{L}^{\ast}$
The impact of the threshold radius $r_{L}^{\ast}$ on the proposed method is examined based on the test-problem results. The threshold radius $r_{L}^{\ast}$ is an important design parameter because it determines the set of effective lobes and affects potential transfer paths. To this end, the test problem with the targeting strategy is solved for $r_{L}^{\ast}=0.003$. Note that the threshold weight $w^{\ast}$ is set to $110$ m/s, as no feasible transfer exists when $w^{\ast}=100$ m/s. Figure 22 shows swarm charts of the total transfer costs for $r_{L}^{\ast}=0.002$ and $0.003$. As expected, a larger $r_{L}^{\ast}$ results in a smaller set of effective lobes in the graph, and fewer transfer trajectories are available. Figure 23 indicates the optimal transfer path when $r_{L}^{\ast}=0.003$. This figure demonstrates that fewer lobes are available, particularly around the 77:22 unstable resonant orbit, resulting in an increase in the total $\Delta V$ by about $43$ m/s. These results support the selection of $r_{L}^{\ast}=0.002$ as a representative value for the following sections.
Figure 22: Swarm charts for the transfer costs of all potential paths for $r_{L}^{\ast}=0.002$ and $0.003$.
Figure 23: Optimal transfer path in the periapsis Poincaré map when $r_{L}^{\ast}=0.003$ (Total $\Delta V=196.2929$ m/s).